# StateFlow: Enhancing LLM Task-Solving through State-Driven Workflows

### v4

@@ B.1 Experiment Details
We refer to Table  11, 12, 13 for the prompts we used. The type of input task, as a prior knowledge, is used in all methods. ReAct and ALFChat use different few-shot examples for different types of tasks. Similarly, we have different planning examples for different tasks, as shown in Table 11. In Process, we also have three different prompts corresponding to heat, cool, and clean. In Figure 8, we illustrate activated states for different tasks. As discussed in the main paper, we identify the object of interest by calling the same LLM at the beginning of the task. In Pick, only when we detect a string match of “You pick up A” (A is the object needed), we would transit the next state. Similarly, in FindLamp, we transit to UseLamp only if a string “desklamp” is matched. In Process, we match the pattern “You heat/cool/clean” to proceed. Note that this feedback is from the environment. Finally at state Put and UseLamp, we transit to End only if the task succeeds or fails. A task is considered success if we receive “Done=True” from environment, and considered fail if the same response is generated by the LLM for three consecutive rounds (following AutoGen). In this experiment, the environment feedback “Done” is available to all methods to terminate the process upon success. Please see Figure 11 for an example.
@@ B.3 Additional Experiments
Table 10: Performance and cost on GPT-3.5-instruct. Here we report success rate of only one attempt. The performances of both methods drop compared to result with GPT-3.5-Turbo, on which ReAct achieves 55.5% and StateFlow achieves 83.3% in overall success rate.
@@
Interact with a MySQL Database system using SQL queries to answer a question.

### v5

@@ B.1 Experiment Details
We refer to Table  12, 13, 14 for the prompts we used. The type of input task, as a prior knowledge, is used in all methods. ReAct and ALFChat use different few-shot examples for different types of tasks. Similarly, we have different planning examples for different tasks, as shown in Table 12. In Process, we also have three different prompts corresponding to heat, cool, and clean. In Figure 8, we illustrate activated states for different tasks. As discussed in the main paper, we identify the object of interest by calling the same LLM at the beginning of the task. In Pick, only when we detect a string match of “You pick up A” (A is the object needed), we would transit the next state. Similarly, in FindLamp, we transit to UseLamp only if a string “desklamp” is matched. In Process, we match the pattern “You heat/cool/clean” to proceed. Note that this feedback is from the environment. Finally at state Put and UseLamp, we transit to End only if the task succeeds or fails. A task is considered success if we receive “Done=True” from environment, and considered fail if the same response is generated by the LLM for three consecutive rounds (following AutoGen). In this experiment, the environment feedback “Done” is available to all methods to terminate the process upon success. Please see Figure 11 for an example.
@@ B.3 Additional Experiments
Addition results with open-source models We further test StateFlow and ALFChat with 3 agents on 3 open-source models: Mistral-7b, Llama-8b and Llama3-70b. We observe that StateFlow outperforms 3-agent ALFChat with 3 agents across the open-source models. It is worth noting that Mistral-7b struggles to complete the task using both methods, suggesting the model itself is not equipped to handle tasks within ALFWorld. However, Llama3-70b with StateFlow can hit a 94% accuracy, surpassing the performance of GPT-3.5 Turbo.
ALFChat (3 agents) StateFlow
Acc (%) Cost (\$) Acc (%) Cost (\$)
Mistral-7b 1.5 0.708 6 2.03
Llama3-8b 50.7 0.52 62.7 0.38
Llama3-70b 83.6 2.53 94 1.83
Table 10: Accuracy and cost comparison on ALFWorld with different open-source models. For this experiment, we compare with the strongest baseline and run the test one time.
@@
Table 11: Performance and cost on ALFWorld with GPT-3.5-instruct. Here we report success rate of only one attempt. The performances of both methods drop compared to result with GPT-3.5-Turbo, on which ReAct achieves 55.5% and StateFlow achieves 83.3% in overall success rate.
@@
Interact with a self.setting system using self.language queries to answer a question.